OpenStudy (anonymous):

(y/(x-9y)=6x^2-4) find the slope of the tangent line to the curve using implicit differentiation the points (1,-3/-26)

OpenStudy (anonymous):

You can do this.

OpenStudy (anonymous):

i used a quotient rule for (y/(x-9y) but i ran it to a problem

OpenStudy (anonymous):

what was it?

OpenStudy (anonymous):

\[ \left(\frac{y}{x-9y}\right)'=\frac{y'(x-9y)-y(1-9y')}{(x-9y)^2} \]

OpenStudy (anonymous):

Is this what you did?

OpenStudy (anonymous):

this is wat i had (x-9y)(y')-y(x-9y')/(x-9y)^2 = 6x

OpenStudy (anonymous):

\[ (x-9y)'=1-9y' \]

OpenStudy (anonymous):

\[ (6x^2-4)'=12x \]

OpenStudy (anonymous):

so how can i isolate the y'?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

help me with this one sin(xy)+x+y^{5} = c . Use implicit differentiation to find the derivative of the family of curves

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Well, use chain rul

OpenStudy (anonymous):

\[ (xy)'=y+xy' \]

OpenStudy (anonymous):

So\[ \cos(xy)(y+xy')+1+5y^4y' \]